OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (texaschic101):

first you need to find the slope by using the slope formula.. slope(m) = (y2 - y1) / (x2 - x1) (7,-3) x1 = 7 and y1 = -3 (4,-8) x2 = 4 and y2 = -8 now we sub slope(m) = (-8 - (-3) / (4 - 7) slope(m) = (-8 + 3) / -3 slope(m) = -5/-3 which equals 5/3 so our slope is 5/3 now you can either use point slope formula (y - y1 = m(x - x1) or you can use slope intercept form (y = mx + b). I prefer slope intercept form unless the problem specifically asks for point slope form, but that is just me. y = mx + b slope(m) = 5/3 you can use either of your sets of points (7,-3) x = 7 and y = -3 now we sub -3 = 5/3(7) + b -3 = 35/3 + b -3 - 35/3 = b -9/3 - 35/3 = b -44/3 = b so your equation is : y = 5/3x - 44/3 but we need it in standard form y = 5/3x - 44/3 --- multiply entire equation by 3 3y = 5x - 44 -- subtract 5x from both sides -5x + 3y = -44 -- multiply by -1 to make x positive 5x - 3y = 44 ==> standard form any questions ?